# Data-dependent and Oracle Bounds on Forgetting in Continual Learning

### v2

@@ Appendix A Appendix - proofs
Starting from Lemma A.1, we can choose $f(z)=\lambda_{t}(\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t}))$, giving us
@@
Corollary A.4.
Restatement of Corollary 3.2: If $\ell\in[0,K]$, for any fixed $S_{s},Q_{s},\lambda_{t}>0$, the following applies uniformly for all posteriors $Q_{s:t}$ with probability at least $1-\delta$ over the choice of $S_{t}$,
@@
Similarly to Lemma A.3 we can apply Fubini’s theorem, and since $\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})$ is independent of $S_{t}$ we can re-order the expectations and use Hoeffding’s lemma on $e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{t},S_{t})}$ if $\ell\in[0,K]$. More formally,
@@
$=\mathbb{E}_{Q_{s}}\left[\mathbb{E}_{S_{t}}\left[e^{\lambda_{t}\Delta\hat{\mathcal{L}}(h,\mathcal{D}_{t},S_{t})}\right]e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}\right]\leq\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}^{2}K^{2}/8m_{t}}e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}\right]$
@@
Theorem A.5.
Restatement of Theorem 4.1: For any $Q_{s},S_{s},\lambda_{t}>0$, if $\ell\in[0,K]$,
@@
Starting from Lemma A.1, we know that
@@
In particular, for $\hat{\rho}_{\lambda}(z)\propto\pi(z)e^{-\lambda_{t}f(z)}$, this is an equality (from Donsker and Varadhan’s [1975] variational lemma). From this, we know that
@@
$=\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
$=\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
@@
Proof of accompanying Corollary:
@@
From our assumption $0\leq\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\epsilon_{s,t}$ we also have $0\leq|\mathcal{L}(Q_{s},\mathcal{D}_{s})-\mathcal{L}(Q_{s},\mathcal{D}_{t})|\leq\epsilon_{s,t}$ since
$\mathbb{E}_{h\sim Q_{s}}\left[\Delta\mathcal{L}(h,s,t)\right]\leq\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t}),\quad\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\max_{h}\left[\Delta\mathcal{L}(h,s,t)\right].$
Looking at the negative transfer, we see
$\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}F(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\left[\mathcal{L}(Q_{s},\mathcal{D}_{t})-\mathcal{L}(Q_{s},\mathcal{D}_{s})\right]+\frac{\lambda_{t}K^{2}}{8m_{t}}+\epsilon_{s,t}$
and similarly from our assumption, we have
$\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}F(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+2\epsilon_{s,t}$
Corollary A.6.
Restatement of Corollary 4.3: If $\ell\in[0,K]$, for any given $S_{s}\sim\mathcal{D}_{s},Q_{s},\lambda_{t}>0$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}&\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\mathcal{L}(h,\mathcal{D}_{s})\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}\end{split}$ (19)
Proof.
We can derive such a bound using (7) by setting the optimal posterior:
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})&\leq-\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{t})}\right]\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]\end{split}$
This is the same as writing
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{t})}\right]}$
Since $e^{k}\geq 0$ for all $k\in\mathbb{R}$, we can apply the log-sum inequality:
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\log\frac{e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}}{e^{\lambda_{t}(-\mathcal{L}(h,\mathcal{D}_{t}))}}\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}$
Corollary A.7.
Appendix-only Corollary: If $\ell\in[0,K]$, for any given $Q_{s},\lambda_{t}>0,S_{s}\sim\mathcal{D}_{s}$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},&\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}\\ &+\frac{\sqrt{\mathbb{E}_{Q_{s}}\left[e^{2\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}}{\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}\cdot\left(\sqrt{\mathrm{Var}_{Q_{s}}(\mathcal{L}(h,\mathcal{D}_{s}))}+\mathcal{L}(Q_{s},\mathcal{D}_{s})\right)\end{split}$ (20)
Proof.
Starting from (4.3), we can apply the Cauchy-Schwartz theorem on the expectation and get
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\sqrt{\mathbb{E}_{Q_{s}}\left[e^{2\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]\mathbb{E}_{Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})^{2}\right]}}{\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]};$
$=\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\sqrt{\mathbb{E}_{Q_{s}}\left[e^{2\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}}{\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\right]}\sqrt{\mathbb{E}_{Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})^{2}\right]}$ (21)
And using the definition of variance and the triangle inequality (since both terms are non-negative), we get (20).
Theorem A.8.
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\mathcal{L}(P,\mathcal{D}_{i})\end{split}$ (22)
Proof.
Starting from (3.3), suppose we assume that for each task we apply an empirical Gibbs learner, meaning
@@
We can unravel the expectations and arrive at:
@@
As in the paper, we mark
@@
And if this covariance is non-negative, we can set it at 00 and still retain a valid upper bound, giving us the forgetting bound via Jensen’s inequality.
@@
and we can further simplify this by taking a slightly looser condition:

### v3

@@ Appendix B Appendix - proofs
Starting from Lemma B.1, we can choose $f(z)=\lambda_{t}(\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t}))$, giving us
@@
Corollary B.4.
If $\ell\in[0,K]$, for any fixed $S_{s},Q_{s},\lambda_{t}>0$, the following applies uniformly for all posteriors $Q_{s:t}$ with probability at least $1-\delta$ over the choice of $S_{t}$,
@@
Similarly to Lemma B.3 we can apply Fubini’s theorem, and since $\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})$ is independent of $S_{t}$ we can re-order the expectations and use Hoeffding’s lemma on $e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{t},S_{t})}$ if $\ell\in[0,K]$. More formally,
@@
$\mathbb{E}_{Q_{s}}\left[\mathbb{E}_{S_{t}}\left[e^{\lambda_{t}\Delta\hat{\mathcal{L}}(h,\mathcal{D}_{t},S_{t})}\right]e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}\right]\leq\mathbb{E}_{Q_{s}}\left[e^{\lambda_{t}^{2}K^{2}/8m_{t}}e^{\lambda_{t}\Delta\mathcal{L}(h,\mathcal{D}_{s},\mathcal{D}_{t})}\right]$
@@
Corollary B.5.
Restatement of Corollary 4.1: For any $Q_{s},S_{s},\lambda_{t}>0$, if $\ell\in[0,K]$,
@@
Starting from Lemma B.1, we know that
@@
In particular, for $\hat{\rho}_{\lambda}(z)\propto\pi(z)e^{-\lambda_{t}f(z)}$, this is an equality (from Donsker & Varadhan’s [1975] variational lemma). From this, we know that
@@
$\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
$\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}))}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{t}))}\right]$
@@
Theorem B.6.
Appendix-only Theorem: For any $\lambda_{T}>0$, assuming all $Q_{1:j}$ are empirical Gibbs posteriors, $\ell\in[0,K]$, the following applies a.s. in expectation over $S_{j}\sim\mathcal{D}_{j}$,
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{i}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}\end{split}$ (15)
@@
By applying Theorem 3.1 with $T$ tasks, we have
$\begin{split}\mathcal{L}&(Q_{1:T},\mathcal{D}_{t})\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})+\frac{1}{\lambda_{T}}\log\mathbb{E}_{h\sim Q_{1:T-1}}\left[e^{\lambda_{T}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{T}))}\right]\end{split}$ (16)
, suppose we assume that for each task we apply an empirical Gibbs learner, meaning
@@
By using the definition of the Gibbs posterior, we can unravel the expectations:
$\begin{split}\forall i\in[1,T-1]:\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{1}{\lambda_{T}}\log\frac{\mathbb{E}_{h\sim\hat{Q}^{\lambda_{T-2}}_{1:T-2}}\left[e^{\lambda_{T}(\mathcal{L}(h,\mathcal{D}_{i})-\hat{\mathcal{L}}(h,S_{T}))-\lambda_{T-1}\mathcal{L}(h,\mathcal{D}_{T-1})}\right]/Z_{T-1}}{\mathbb{E}_{h\sim\hat{Q}^{\lambda_{T-2}}_{1:T-2}}\left[e^{-\lambda_{T-1}\mathcal{L}(h,\mathcal{D}_{T-1})-\lambda_{T}\hat{\mathcal{L}}(h,S_{T})}\right]/Z_{T-1}}\end{split}$
Setting $\lambda_{T-1}=\lambda_{T}$ and repeating this process, we arrive at:
@@
Theorem B.7.
Restatement of Theorem 4.3 For any $\lambda_{T}>0$, assuming all $Q_{1:j}$ are empirical Gibbs posteriors, and that Assumption 2 holds, the following holds a.s. in expectation over $S_{j}\sim\mathcal{D}_{j}$, $\forall j\in[T]$
$\begin{split}\lim_{m,T\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S_{1},\ldots,S_{T}}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\lim_{m,T\rightarrow\infty}\frac{1}{T-1}\sum_{i=1}^{T-1}\mathcal{L}(h^{*}_{1:T/i},\mathcal{D}_{i}).\end{split}$ (17)
Proof.
Starting from (15), we have (assuming all training sets are of size $m$ and $\lambda_{i}=\lambda_{T}$) $\forall i\in[1,T-1]$
$\begin{split}\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m}+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{i}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}(\hat{\mathcal{L}}(h,S_{j})-\mathcal{L}(h,\mathcal{D}_{j})+\mathcal{L}(h,\mathcal{D}_{j}))}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}(\hat{\mathcal{L}}(h,S_{j})-\mathcal{L}(h,\mathcal{D}_{j})+\mathcal{L}(h,\mathcal{D}_{j}))}\right]}\end{split}$
Taking the expectation on all training sets (marked $\mathbb{E}_{S}$ for brevity), we can apply Jensen’s inequality as well as Fubini’s Theorem to obtain $\forall i\in[1,T-1]$
$\begin{split}\mathbb{E}_{S}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{1}{\lambda_{T}}\log\mathbb{E}_{S}\frac{\mathbb{E}_{h\sim P}\left[\mathbb{E}_{S}\left[e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}(\hat{\mathcal{L}}(h,S_{j})-\mathcal{L}(h,\mathcal{D}_{j}))}\right]e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}(\hat{\mathcal{L}}(h,S_{j})-\mathcal{L}(h,\mathcal{D}_{j})+\mathcal{L}(h,\mathcal{D}_{j}))}\right]}\end{split}$
Via a similar application of Hoeffding’s lemma as in Lemma B.3, we have (for $m\rightarrow\infty$)
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{\lambda_{T}K^{2}T}{8m}+\frac{1}{\lambda_{T}}\log\mathbb{E}_{S}\frac{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}(\hat{\mathcal{L}}(h,S_{j})-\mathcal{L}(h,\mathcal{D}_{j})+\mathcal{L}(h,\mathcal{D}_{j}))}\right]}\end{split}$
If we have (for $m\rightarrow\infty$)
$cov_{P}\left(e^{\lambda_{T}\sum_{j=1}^{T}\mathcal{L}(h,\mathcal{D}_{j})},e^{-\lambda_{T}\sum_{j=1}^{T}\hat{\mathcal{L}}(h,S_{j})}\right)\leq 0$
we also have
$cov_{P}\left(e^{-\lambda_{T}\sum_{j=1}^{T}\mathcal{L}(h,\mathcal{D}_{j})},e^{-\lambda_{T}\sum_{j=1}^{T}\hat{\mathcal{L}}(h,S_{j})+\mathcal{L}(h,\mathcal{D}_{j}))}\right)\geq 0$
and therefore (for $m\rightarrow\infty$)
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\ldots+\frac{1}{\lambda_{T}}\log\mathbb{E}_{S}\frac{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}\hat{\mathcal{L}}(h,S_{j})-\mathcal{L}(h,\mathcal{D}_{j})}\right]\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}\end{split}$
$\begin{split}=\ldots+\frac{1}{\lambda_{T}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}\mathbb{E}_{S}\frac{1}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}\hat{\mathcal{L}}(h,S_{j})-\mathcal{L}(h,\mathcal{D}_{j})}\right]}\end{split}$
Via Jensen’s inequality
$\begin{split}\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{\lambda_{T}K^{2}T}{8m}+\frac{1}{\lambda_{T}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}\mathbb{E}_{S}\mathbb{E}_{h\sim P}\left[e^{\lambda_{T}\sum_{j=1}^{T}\hat{\mathcal{L}}(h,S_{j})+\mathcal{L}(h,\mathcal{D}_{j})}\right]\end{split}$
Via Fubini’s theorem and Hoeffding’s lemma, similarly to Lemma B.3,
$\begin{split}\forall i\in[1,T-1]:\mathbb{E}_{S}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\frac{\lambda_{T}K^{2}T}{4m}+\frac{1}{\lambda_{T}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1,j\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\sum_{j=1}^{T}\mathcal{L}(h,\mathcal{D}_{j})}\right]}\end{split}$
Averaging over tasks, we have
$\begin{split}\frac{1}{T-1}\mathbb{E}_{S}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m}+\frac{\lambda_{T}K^{2}T}{4m}+\\ &\frac{1}{\lambda_{T}(T-1)}\mathbb{E}_{S}\sum_{i=1}^{T-1}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}(T-1)\sum_{j=1,j\neq i}^{T}\frac{\mathcal{L}(h,\mathcal{D}_{j})}{(T-1)}}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\mathcal{L}(h,\mathcal{D}_{i})}e^{-\lambda_{T}(T-1)\sum_{j=1,j\neq i}^{T}\frac{\mathcal{L}(h,\mathcal{D}_{j})}{(T-1)}}\right]}\end{split}$
Under Assumption 2, we can apply Laplace’s method (Shun & McCullagh, 1995) on both numerator and denominator assuming $\lambda_{T}(T-1)\rightarrow\infty$. Taking $\lambda_{T}=\sqrt{m}/T$ and considering the limit $m,T\rightarrow\infty$ we therefore have
$\begin{split}&\lim_{m,T\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\sqrt{m}/T}_{1:T},\mathcal{D}_{i})\leq\lim_{m,T\rightarrow\infty}\frac{T}{\sqrt{m}(T-1)}\mathbb{E}_{S}\sum_{i=1}^{T-1}\log A_{i}\end{split}$
where
$A_{i}=\frac{C(m,T,d,h^{*}_{1:T/i})\cdot\left[e^{-\sqrt{m}(T-1)/T\sum_{j=1,j\neq i}^{T}\frac{\mathcal{L}(h^{*}_{1:T/i},\mathcal{D}_{j})}{(T-1)}}\right]\left(1+O(\frac{T}{\sqrt{m}T})+\ldots\right)}{C(m,T,d,h^{*}_{1:T/i})\cdot\left[e^{-\sqrt{m}/T\mathcal{L}(h^{*}_{1:T/i},\mathcal{D}_{i})}e^{-\sqrt{m}(T-1)/T\sum_{j=1,j\neq i}^{T}\frac{\mathcal{L}(h^{*}_{1:T/i},\mathcal{D}_{j})}{(T-1)}}\right]\left(1+O(\frac{T}{\sqrt{m}T})+\ldots\right)}.$
Since most of the elements in the numerator and denominator are shared, we have
$\begin{split}&\lim_{m,T\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\sqrt{m}/T}_{1:T},\mathcal{D}_{i})\leq\lim_{m,T\rightarrow\infty}\frac{T}{\sqrt{m}(T-1)}\mathbb{E}_{S}\sum_{i=1}^{T-1}\log\frac{1}{e^{-\sqrt{m}/T\hat{\mathcal{L}}(h^{*}_{1:T/i},S_{i})}}\end{split}$
A rearrangement of terms completes the proof. ∎
Theorem B.8.
$\begin{split}\forall i\in[1,T]:\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\frac{\lambda_{T}K^{2}}{8m_{i}}+\mathcal{L}(P,\mathcal{D}_{i})\end{split}$ (18)
Proof.
Starting from (15), we mark
@@
And if this covariance is non-negative, we can set it at 0 and still retain a valid upper bound, giving us the forgetting bound via Jensen’s inequality.
@@
We note that since $\ell\in[0,K]$, we have $0<\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]\leq 1.$
We can further simplify this by taking a slightly looser condition: